OpenStudy (anonymous):

Please help / Fan and Medal. I'm dying here ): Susi and Janet have been solving systems of equations with one polynomial function of degree two or higher and one linear function. Janet says there must always be one solutions, and Susi says there will always be two solutions. Using complete sentences, explain how Susi can be correct, how Janet can be correct, and how they both can be wrong.

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

so image we have the function f(x)=ax^2+bx+c

myininaya (myininaya):

you know the discriminant is b^2-4ac when the discriminant is 0 how many real solutions do we have? when the discriminant is positive how many real solutions do we have? when the discriminant is negative how many real solutions do we have?

OpenStudy (anonymous):

Could you right out the equation, I forgot what it looked like and typing the question confuses me

myininaya (myininaya):

what equation are you asking about?

myininaya (myininaya):

I was going to first have you consider functions in the form y=ax^2+bx+c and then use that same knowledge to provide a conjecture about functions in the form y=ax^3+bx^2+cx+d. Then we could say something about polynomial functions of degree n.

myininaya (myininaya):

for example, for the 2nd degree polynomial pretend we have f(x)=x^2 <--how many real solutions does this one have? f(x)=x^2+1 <--how many real solutions does this one have? f(x)=x^2-1 <--how many real solutions does this one have?

OpenStudy (anonymous):

the first one, zero. the second, one. and the third one, two. am I correct? @myininaya

myininaya (myininaya):

yes so what does that mean f(x)=ax^2+bx+c can have at most 2 real solutions since we seen it could have either 0, 1 , or 2 solutions.

myininaya (myininaya):

so if a f(x)=ax^2+bx+c can have at most 2 real solutions then f(x)=ax^3+bx^2+cx+d can have at most what real solutions?

OpenStudy (anonymous):

two solutions?

OpenStudy (anonymous):

THREE! sorry

myininaya (myininaya):

well a second degree polynomial has a most 2 real solutions so third degree polynomial would have most how many real solutions?

myininaya (myininaya):

yes! :)

myininaya (myininaya):

at most means it doesn't have to be exactly 3 it could be 2 1 or even 0

OpenStudy (anonymous):

So how exactly could I answer this question? Im not exactly good at this ): I know that if the function touched the line one time, theres one solution. I know that if It hits it twice, theres two solutions. if the parabola doesn't touch the line at all, there is no solution... but how exactly can I answer this question? help?

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

that sounds like a good start for the 2nd degree poly you can expand what you said for a 3rd degree on (except don't call it a parabola because it isn't )

myininaya (myininaya):

I would also provide at least the parabola examples too.